### v4

# Superconductivity in twisted transition metal dichalcogenide homobilayers

Abstract
Recently, robust superconductivity has been independently reported in twisted WSe2 bilayers by two separate groups [Y. Xia et al., arXiv:2405.14784; Y. Guo et al., Nature 637, 839–845 (2025)]. In light of this, we explore the possibility of a universal superconducting pairing mechanism in twisted WSe2 bilayers. Using a continuum band structure and a phenomenological boson-mediated effective electron-electron attraction, we find that intervalley intralayer pairing predominates over interlayer pairing. Notably, despite very different experimental conditions, both twisted WSe2 samples exhibit a comparable effective attraction strength for superconductivity. This consistency in the two experiments suggests that the dominant pairing glue is likely independent of the twist angle and layer polarization, pointing to a universal underlying boson-induced pairing mechanism. We speculate on the possible mechanisms in the discussion.
Introduction.— The emergence of moiré superlattice in twisted transition metal dichalcogenide (TMD) homobilayers has led to the observation of a rich set of strongly correlated phases Tang et al. (2020); Regan et al. (2020); Xu et al. (2020); Wang et al. (2020a); Li et al. (2021a); Ghiotto et al. (2021); Li et al. (2021b); Foutty et al. (2024); Cai et al. (2023); Zeng et al. (2023); Park et al. (2023); Xu et al. (2023), from correlated insulators at integer fillings to fractional Chern insulators in the absence of an external magnetic field. While the phases driven by strong electron-electron interaction manifest in the moiré TMD bilayers, robust observable superconductivity (SC) at dilute doping densities–associated with a few carriers per moiré unit cell–has remained elusive in most moiré TMD systems Wang et al. (2020b). On the other hand, both moiré and moiréless graphene multilayers have demonstrated reproducible SC with $T_{c}$ ranging from $20$ mK to $3$ K Cao et al. (2018); Yankowitz et al. (2019); Lu et al. (2019); Park et al. (2021); Hao et al. (2021); Cao et al. (2021); Oh et al. (2021); Zhou et al. (2021, 2022); Su et al. (2023); Zhang et al. (2023); Holleis et al. ; Li et al. (2024), indicating that SC is a generic feature in graphene-based multilayers. This difference between moiré TMD and moiré graphene systems is an important unsolved puzzle in condensed matter and material physics.
Recently, two groups have independently discovered robust SC in twisted WSe2 bilayers (tWSe2). One group reported SC at a filling factor $\nu_{\rm SC}=-1$ with a twist angle of $\theta=3.65^{\circ}$ and a small displacement field Xia et al. . The critical temperature was estimated to be $0.22$ K. The other group observed SC around $\nu_{\rm SC}=-1.1$ with a twist angle of $\theta=5^{\circ}$ and a large displacement field Guo et al. . The critical temperature was approximately $0.426$ K. We will refer to these two experimental samples as Sample A Xia et al. and Sample B Guo et al. , respectively, throughout our paper.
In this Letter, we investigate SC in tWSe2 using a realistic continuum band structure model and a phenomenological boson-mediated BCS model without assuming a specific pairing mechanism. We demonstrate the dominance of intervalley intralayer pairing with an order parameter consistent with a mixture of $s$ and $f$ waves Hsu et al. (2021); Chou et al. (2022a); Akbar et al. ; pai . Our estimates of the effective coupling constants of SC in the two experiments show very similar values, suggesting a universal pairing mechanism in tWSe2 (i.e., the same bosonic glue). The possible microscopic mechanisms and the importance of the in-plane magnetic field response are also discussed. Our work represents the first systematic investigation of the SC phenomenology in tWSe2 systems, paving the way for future exploration of SC in moiré TMD systems.
Band Structure Model.— Our phenomenological superconducting mean-field theory is based on the low-energy continuum model of tWSe2 Wu et al. (2019a), of which the valley-projected Hamiltonian is
$\mathcal{H}=\begin{pmatrix}h(\boldsymbol{q}_{1})+U_{1}(\boldsymbol{r})+\frac{V_{z}}{2}&T(\boldsymbol{r})\\ T^{\dagger}(\boldsymbol{r})&h(\boldsymbol{q}_{2})+U_{2}(\boldsymbol{r})-\frac{V_{z}}{2}\end{pmatrix},$ (1)
where $h(\boldsymbol{q})=-\hbar^{2}q^{2}/2m^{*}$ is the effective mass approximation of the valence band edge at the Dirac point $\boldsymbol{\kappa}_{l}$ of layer $l=1,2$. $\boldsymbol{q}_{l}=\boldsymbol{k}-\boldsymbol{\kappa}_{l}$ is the momentum measured from the Dirac points, and $V_{z}$ is the interlayer energy difference due to the displacement field. The layer-dependent moiré potential $U_{l}(\boldsymbol{r})$ and the interlayer tunneling $T(\boldsymbol{r})$ are spatially periodic with moiré periodicity,
$U_{1,2}(\boldsymbol{r})=2V\sum\limits_{j=1,3,5}\cos(\boldsymbol{G}_{j}\cdot\boldsymbol{r}\mp\psi),$ (2)
$T(\boldsymbol{r})=w(1+e^{-i\boldsymbol{G}_{2}\cdot\boldsymbol{r}}+e^{-i\boldsymbol{G}_{3}\cdot\boldsymbol{r}}).$ (3)
$\boldsymbol{G}_{1}=(1,0)b_{\rm M}$, $\boldsymbol{G}_{j}=\mathcal{R}_{(j-1)\pi/3}\boldsymbol{G}_{1}$ are the first-shell moiré reciprocal lattice vectors, with $b_{\rm M}=4\pi\theta/\sqrt{3}a_{0}$ as the length of moiré primitive reciprocal lattice vector. In the calculations throughout this paper, we use the continuum model parameters fitted by large-scale DFT calculations Devakul et al. (2021): $V=9$ meV, $\psi=128^{\circ}$, $w=18$ meV, lattice constant $a_{0}=3.317\text{ \AA}$, and the effective mass $m^{*}=0.43m_{e}$ with $m_{e}$ being the electron mass.
Figure 1: Band properties of Sample A: $\theta=3.65^{\circ}$, $V_{z}=2.1$ meV and the filling factor of SC $\nu_{\rm SC}=-1$. (a) The band structure along high-symmetry lines and the energy contour (inset) of the first moiré valence band. The dashed horizontal line and the dashed contour in the inset represent the chemical potential and the Fermi surface at $\nu_{\rm SC}$, respectively. The color of the first moiré band indicates the layer polarization $l_{z}$. (b) The DOS for $\nu\in[-2,0]$, taking into account both valleys, i.e., the filling factor $\nu$ denotes the number of electrons per moiré unit cell. The VHS appears around $\nu=-0.77$ and is $\sim 1.5$ meV away from the chemical potential at $\nu_{\rm SC}$ (inset). (c) The layer polarization distribution in the first MBZ. The white dashed line outlines the Fermi surface contour at $\nu_{\rm SC}$. (d-e) The layer-projected densities, $\rho_{1}$ and $\rho_{2}$, in real space, forming two triangular lattices centered on XM or MX local stackings. $\rho_{1}$ and $\rho_{2}$ create an effective honeycomb lattice with a small onsite energy difference. The white hexagon marks the moiré unit cell.
Figure 2: Band properties of Sample B: $\theta=5^{\circ}$, $V_{z}=43.75$ meV and the filling factor of SC $\nu_{\rm SC}=-1.1$. (a) The band structure along high-symmetry lines and the energy contour (inset) of the first moiré valence band. The dashed horizontal line and the dashed contour in the inset represent the chemical potential and the Fermi surface at $\nu_{\rm SC}$, respectively. (b) The DOS for $\nu\in[-2,0]$, showing that the VHS is near $\nu_{\rm SC}$ and less than $0.5$ meV from the chemical potential at $\nu_{\rm SC}$ (inset). (c) The layer polarization distribution in the first MBZ. The white dashed line outlines the Fermi surface contour at $\nu_{\rm SC}$. (d-e) The layer-projected density distributions in real space.
Figures 1 and 2 show the moiré band structures, density of states (DOS) and layer polarization $l_{z}$ in momentum and real space for Samples A and B. The layer polarization operator $\hat{l}_{z}$ is represented by the Pauli matrix acting on the layer subspace. In Sample A, the layer is weakly polarized due to the small displacement field, corresponding to an interlayer energy difference of approximately $2.1$ meV Xia et al. ; Wef . In momentum space, as shown in Fig. 1(c), the wave function is localized on each layer around its Dirac point. In real space, depicted in Fig. 1(d-e), the layer projected density forms two triangular lattices centered on XM or MX local stackings, creating an effective honeycomb lattice with a small onsite energy difference. The DOS plot in Fig. 1(b) shows a Van Hove singularity (VHS) around $\nu=-0.77$. The SC observed at $\nu_{\rm SC}=-1$ is only $\sim 1.5$ meV away from the VHS (inset of Fig. 1(b)). In our notation, the filling factor $\nu$ represents the number of electrons per moiré unit cell. In Sample B, a large displacement field induces an interlayer energy difference of $\sim 43.75$ meV Guo et al. ; Wea , resulting in strong layer polarization (Fig. 2(c-e)). The VHS is located at $\nu=-1.14$ (Fig. 2(b)), which is less than $0.5$ meV away from the chemical potential at the filling factor $\nu_{\rm SC}=-1.1$, where SC was observed. This proximity is conducive to pairing instability.
Despite the significant differences in band structure, layer polarization, and effective lattice model between Samples A and B, we explore the possibility that the same underlying mechanism is responsible for SC in both cases in the next section.
Superconductivity.— Closely connected to the two experiments, we study the effective pairing strength in Samples A and B using a phenomenological BCS theory without assuming a specific pairing mechanism. We will discuss the possible pairing glues, e.g., phonons and magnons, at the end of this Letter. Given that no evidence of time-reversal symmetry (TRS) breaking was observed near the superconducting phase in either sample, we consider only the intervalley pairing that preserves TRS. In the following part, we focus on intralayer pairing, as detailed in Supplementary Information (SI) A, the interlayer pairing is significantly weaker than intralayer pairing by a factor of five to ten. The dominance of the intralayer pairing is a direct consequence of the layer polarization patterns as shown in Figs. 1(c) and 2(c).
The effective electron-electron attraction mediated by intralayer bosons is given by
$\begin{split}H_{\rm att}=-\frac{g}{2A}\sum\limits_{l,\boldsymbol{q}}\hat{n}_{l}(\boldsymbol{q})\hat{n}_{l}(-\boldsymbol{q}),\end{split}$ (4)
where $A$ is the system area, and we approximate the intervalley intralayer pairing strength by a tunable static and momentum independent potential $g$. The density operator, $\hat{n}_{l}(\boldsymbol{q})$, is defined as
$\hat{n}_{l}(-\boldsymbol{q})=\sum\limits_{\tau,\boldsymbol{k},\boldsymbol{G}}\psi^{\dagger}_{\tau,l}(\boldsymbol{k}+\boldsymbol{G})\psi_{\tau,l}(\boldsymbol{k}+\boldsymbol{G}-\boldsymbol{q}),$ (5)
where $\tau=\pm$ is the valley index, $l=1,2$ represents the layer degree of freedom and $\boldsymbol{k}$ is in the first moiré Brillouin zone (MBZ). Thus, the effective attraction Eq. (4) is
$\begin{split}H_{\rm att}=-\frac{g}{A}\sum\limits_{\begin{subarray}{c}l,\boldsymbol{k},\boldsymbol{k}^{\prime}\\ \boldsymbol{G},\boldsymbol{G}^{\prime}\end{subarray}}&\psi^{\dagger}_{+,l}(\boldsymbol{k}+\boldsymbol{G})\psi^{\dagger}_{-,l}(-\boldsymbol{k}-\boldsymbol{G})\\ &\psi_{-,l}(-\boldsymbol{k}^{\prime}-\boldsymbol{G}^{\prime})\psi_{+,l}(\boldsymbol{k}^{\prime}+\boldsymbol{G}^{\prime}).\end{split}$ (6)
The factor of 2 in Eq. (4) is canceled by summing over valleys and keeping only the intervalley terms. Projecting to the first moiré valence band of tWSe2, the effective pairing Hamiltonian becomes
$H_{\rm p}=-\frac{1}{A}\sum\limits_{\boldsymbol{k},\boldsymbol{k}^{\prime}}g_{\boldsymbol{k},\boldsymbol{k}^{\prime}}c^{\dagger}_{+}(\boldsymbol{k})c^{\dagger}_{-}(-\boldsymbol{k})c_{-}(-\boldsymbol{k}^{\prime})c_{+}(\boldsymbol{k}^{\prime}),$ (7)
$g_{\boldsymbol{k},\boldsymbol{k}^{\prime}}=g\sum\limits_{l,\boldsymbol{G},\boldsymbol{G}^{\prime}}|z_{+,l,\boldsymbol{G}}(\boldsymbol{k})|^{2}|z_{+,l,\boldsymbol{G}^{\prime}}(\boldsymbol{k}^{\prime})|^{2},$ (8)
where $c^{\dagger}$ ($c$) is the quasiparticle creation (annihilation) operator in the plane-wave expansion,
$c^{\dagger}_{\tau}(\boldsymbol{k})=\sum\limits_{l,\boldsymbol{G}}z_{\tau,l,\boldsymbol{G}}(\boldsymbol{k})\psi^{\dagger}_{\tau,l}(\boldsymbol{k}+\boldsymbol{G}),$ (9)
with $\boldsymbol{G}$ and $\boldsymbol{G}^{\prime}$ being the moiré reciprocal lattice vectors.
For $T\approx T_{c}$, the linearized gap equation in the BCS mean-field approximation is
$\Delta_{\boldsymbol{k}}=\frac{1}{A}\sum\limits_{\boldsymbol{k}^{\prime}}g_{\boldsymbol{k},\boldsymbol{k}^{\prime}}\frac{\tanh(\frac{\varepsilon_{+}(\boldsymbol{k}^{\prime})-\mu}{2k_{\rm B}T})}{2(\varepsilon_{+}(\boldsymbol{k}^{\prime})-\mu)}\Delta_{\boldsymbol{k}^{\prime}},$ (10)
$\varepsilon_{\tau}(\boldsymbol{k})$ is the quasiparticle eigenenergy of the first moiré valence band and $\mu$ is the chemical potential. In obtaining Eq. (8) and (10), we have used the time-reversal symmetry properties
$\varepsilon_{+}(\boldsymbol{k})=\varepsilon_{-}(-\boldsymbol{k}),$ (11)
$z_{+,l,\boldsymbol{G}}(\boldsymbol{k})=z^{*}_{-,l,-\boldsymbol{G}}(-\boldsymbol{k}).$ (12)
Equation (10) can be rewritten in the matrix form:
$\boldsymbol{\Delta}=g\boldsymbol{M}\boldsymbol{\Delta}.$ (13)
At the critical point $T=T_{c}$, Eq. (13) has only one stable solution. Given the electron-boson coupling strength $g$, $T_{c}$ is found by obtaining the largest eigenvalue of $\boldsymbol{M}$ to be $g^{-1}$. Equivalently, for a given temperature $T_{c}$, the critical electron-boson coupling strength $g^{*}$ is the inverse of the maximum eigenvalue of $\boldsymbol{M}$.
[…]
In Fig 3(c-d), $T_{c}$ is shown as a function of $\nu$ for several representative values of $g$. In both samples, $T_{c}$ reaches its maximum at the VHS and remains observable away from the VHS, similar to the situation found in the graphene multilayers Löthman and Black-Schaffer (2017); Chou et al. (2021, 2022b, 2022c). We note that non-adiabatic vertex corrections Cappelluti and Pietronero (1996); Phan and Chubukov (2020), which we ignore, might become important for doping densities very close to VHS. This is an interesting subject for future work. To further explore the dependence of $g^{*}$ and $T_{c}$ on the displacement field, which is a common experimental tuning parameter, we show phase diagrams of $g^{*}$ and $T_{c}$ in SI B for twist angles $\theta=3.65^{\circ}$ and $\theta=5^{\circ}$. In both cases, the minimum $g^{*}$ and maximum $T_{c}$ track the VHS.
[…]
For $T<T_{c}$, the order parameter $\Delta_{\boldsymbol{k}}$ is solved self-consistently,
$\begin{split}\Delta_{\boldsymbol{k}}&=\frac{1}{A}\sum\limits_{\boldsymbol{k}^{\prime}}g_{\boldsymbol{k},\boldsymbol{k}^{\prime}}\frac{\tanh(\frac{\sqrt{\xi^{2}_{\boldsymbol{k}^{\prime}}+|\Delta_{\boldsymbol{k}^{\prime}}|^{2}}}{2k_{\rm B}T})}{2\sqrt{\xi^{2}_{\boldsymbol{k}^{\prime}}+|\Delta_{\boldsymbol{k}^{\prime}}|^{2}}}\Delta_{\boldsymbol{k}^{\prime}},\end{split}$ (14)
where $\xi_{\boldsymbol{k}}=\varepsilon_{+}(\boldsymbol{k})-\mu$. At $T=0$, the hyperbolic tangent term simplifies to 1.
The $k$-space distributions of the order parameters for Samples A and B, calculated at $T=0$, are shown in Fig. 4. Since we consider only intralayer pairing, the symmetry of $\Delta_{\boldsymbol{k}}$ closely matches that of the layer polarization shown in Fig. 1(c) and Fig. 2(c), indicating a mixture of $s$ and $f$ waves Zegrodnik and Biborski (2023); Akbar et al. . In Sample A, $\bar{\Delta}_{k}/k_{B}T_{c}=1.84$, and $\bar{\Delta}_{k}/k_{B}T_{c}=1.66$ in Sample B, where $\bar{\Delta}_{k}$ is the $k$-space average of $\Delta_{k}$. Both ratios are close to the BCS mean-field value of $1.75$.
Figure 4: The order parameter $\Delta_{\boldsymbol{k}}$ for (a) Sample A and (b) Sample B. In (a), $\bar{\Delta}_{k}/k_{B}T_{c}=1.84$ and in (b) $\bar{\Delta}_{k}/k_{B}T_{c}=1.66$, where $\bar{\Delta}_{k}$ is the $k$-space average of $\Delta_{k}$.
Discussion.— We develop a BCS theory for the recently observed SC in tWSe2. We establish that the SC likely arises from the same bosonic glue in two different experiments with different twist angles, displacement fields, and doping levels. The dominant pairing is intervalley intralayer interaction with an $s+f$ order parameter symmetry, and the maximum $T_{c}$ is not far from the VHS. Take the acoustic phonon as an example Wu et al. (2019b); Wu and Das Sarma (2020); Lewandowski et al. (2021); Chou et al. (2021, 2022b, 2022c); Boström et al. for the bosonic glue, the calculated $g^{*}$ is related to the deformation potential by $D=v_{s}\sqrt{g^{*}\rho_{m}}$. Using the mass density $\rho_{m}=6.2\times 10^{-7}$ g/cm2 and sound velocity $v_{s}=3.3\times 10^{5}$ cm/s of monolayer WSe2 Jin et al. (2014), $g^{*}_{\rm SC}$ of Samples A and B in Fig. 3(a-b) correspond to deformation potentials of $7.1$ eV and $6.7$ eV, respectively, which are in the same order of magnitude as $D=3.2$ eV for monolayer WSe2 estimated from previous density functional theory (DFT) calculations Jin et al. (2014). The quantitative extraction of the deformation potential is difficult and DFT estimates are often much smaller than experimental values, a common occurrence in semiconductors.
It is also possible that the SC observed in both samples is mediated by magnons, as suggested by the close proximity of the SC to the correlated state, very likely of an antiferromagnetic order, and the $T_{c}$ that peaks near the phase boundary between these states. Understanding the magnetic order, potentially through a tight-binding model Crépel and Millis , is essential for characterizing the nature of spin fluctuations and the resulting superconducting state. At present, the underlying mechanism of SC remains an unresolved and intriguing question that future experimental and theoretical work should address. Additionally, near a VHS, repulsive interactions may lead to competing orders such as (pseudo)spin density waves, which could influence the SC phase. These competing orders may either coexist with or suppress SC under certain conditions, depending on factors like dielectric constant, doping density and displacement field. A thorough exploration of the interplay between SC and these competing phases remains an open problem and requires further investigation.
Next, we comment on the role of Coulomb interactions. Coulomb interactions may result in correlated states Kim et al. that could preempt SC predicted by our phenomenological BCS theory. The Coulomb repulsion in the Cooper channel is effectively captured by our theory through the effective coupling constant $g$. However, the static intraband screening depends on the density of states, likely resulting in a more drastic change in $T_{c}$ (than our current phenomenological BCS approach) close to VHS Chou et al. (2022b). Microscopic calculations incorporating the frequency-dependent pairing and Coulomb interactions are necessary for a complete quantitative understanding of the underlying SC mechanism in tWSe2 Xia et al. ; Guo et al. , but our current work establishes a bosonic glue to be the likely microscopic mechanism. Moreover, the band renormalization effects due to Coulomb interactions, which we have ignored, may quantitatively modify the band structure used in this Letter. Regarding the single-particle band structure, a more accurate description which utilizes neural networks Zhang et al. (2024) suggest that for large twist angles ($\theta\gtrsim 5^{\circ}$), the topmost moiré valence bands of tWSe2 likely originate from the $\Gamma$-valley, rather than the $K$-valley as assumed in our model. The potential new physics associated with the low-energy bands from the $\Gamma$-valley remains an open question for future investigation.
Finally, we discuss the response to an in-plane magnetic field, which has been an important tool to discern the properties of SC in graphene-based materials Cao et al. (2021); Zhou et al. (2021, 2022); Su et al. (2023); Zhang et al. (2023); Holleis et al. ; Li et al. (2024). In WSe2, the Zeeman effect can be ignored due to large Ising spin-orbit coupling. However, the orbital effect can still be nontrivial as the separation between two WSe2 layers is not small. We find that SC is suppressed by an in-plane magnetic field of a few Teslas because the nesting of intervalley pairing requires $\varepsilon_{+}(\boldsymbol{k})=\varepsilon_{-}(-\boldsymbol{k})$, which is easily violated in the presence of a magnetic field. Additionally, an in-plane magnetic field may influence nearby correlated states; for example, in an antiferromagnetic state, spin fluctuation can be reduced by an applied magnetic field, weakening fluctuation-induced pairing. Systematic investigations along these lines are essential for a complete understanding of SC in tWSe2.
Acknowledgments.— We acknowledge the valuable discussion with Cory R. Dean. Y.-Z. C. and J. Z. also thank Yuting Tan, Yi Huang and Jay D. Sau for useful conversations. This work is supported by the Laboratory for Physical Sciences.